The Dubs division of Fast Company (the parent company) produces wheels for off-road sport vehicles. One-half of Dub's output is sold to the Hoon division of Fast; the remainder is sold to outside customers. Dub's estimated operating profit for the year is shown in the table.

Internal Sales

External Sales

Totals

Sales

$300,000

$400,000

$700,000

Var Mfg.

$160,000

$160,000

$320,000

Var G&A

$40,000

$60,000

$100,000

CM

$100,000

$180,000

$280,000

Fixed Mfg.

$24,000

$32,000

$56,000

Fixed G&A

$36,000

$48,000

$84,000

Op. Profits

$40,000

$100,000

$140,000

Unit Sales

1,000

1,000

2,000

Unless otherwise stated assume the fixed costs given above are allocated costs and unavoidable. Hoon division has an opportunity to purchase 1,000 wheels of the same quality from an outside supplier on a continuing basis for $250.00 per wheel.

To simplify this example, assume the capacity of the Dubs division is 2,000 units and it is producing and selling units as shown in the table. If Dubs could sell all of its units in the external market at the current external selling price, by how much would Fast’s total contribution margin increase? Also assume the Hoon division would acquire its wheels from the external supplier at the stated price.
Answer #1

All amounts are in $

Sale price to outsiders = $400,000/1,000 = $400

Transfer price to Hoon division = $300,000/1,000 = $300

Variable Manufacturing cost per unit = 160,000/1,000 = $160

Variable G&A cost per unit = 60,000/1,000 = $60

(Variable cost of outside sales alone are taken, because we will sell to outsiders totally now)

If Hoon division acquires wheels from outsiders and the Dub's division could sell its total output to outsiders at present price, then

Sale Price = 400

Variable cost per unit = $220 ($160+$60)

Contribution per unit = $180

Total Contribution = 2,000 units x $180 = $360,000

Existing contribution of Dub's division = $280,0000

Increase in Contribution margin for Dub's division = $80,000

Bu we are asked to calculate the increase in contribution margin of the Fast Company as a whole (both divisions combined)

For hoon division, now the varable cost per unit will decrease from $300 (transfer price) to $250 (outside purchase price)

Increase in contribution = 1,000 units x $50 = $50,000

So Total Contribution margin increase for Fast Company is $130,000 ($80,000 + $50,000)

Note :

Contribution margin increase for Dub = $80,000

Contribution margin increase for Hoon = $50,000

Contribution margin increase for Fast Company = $130,000